OpenStudy (anonymous):

A system of equations is shown below: 6x – 2y = 3 (equation 1) 5x + 3y = 4 (equation 2) A student wants to prove that if equation 2 is kept unchanged and equation 1 is replaced with the sum of equation 1 and a multiple of equation 2, the solution to the new system of equations is the same as the solution to the original system of equations. If equation 2 is multiplied by 1, which of the following steps should the student use for the proof?

OpenStudy (anonymous):

Show that the solution to the system of equations 10x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 10x – 2y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 11x – y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations Show that the solution to the system of equations 11x + y = 7 and 5x + 3y = 4 is the same as the solution to the given system of equations

OpenStudy (anonymous):

@kc_kennylau @satellite73

OpenStudy (cgreenwade2000):

Didn't I answer this question already?

OpenStudy (cgreenwade2000):

If you still need help, I'll post it again. All we have to do here is follow the last few steps. They just want us to add Eq1 and some multiple of Eq2. But, they tell you the multiple is 1. So we're really just adding the two equations together to give us a new equation 1.